Hey I'm just so lost, my older sister tried helping me but she can't remember what to do. I would greatly appreciate your help! (I also greatly NEED it) hahaa By the way I'm not sure if I picked the right topic...but I tried. anyways it's bookwork it says:
Sketch each graph using as values of x the integers in the given interval.
y=3^x,[-3,3]

Answer by jim_thompson5910(35256)

To graph on this given interval, we simply plug in values of x that are between -3 and 3 to get corresponding values of y.


These x and y values will pair up to make a series of points that will be plotted on a coordinate system.

From there, you just draw a curve through all the points (the best you can) to graph y = 3^x


So when x = -3, then...

y = 3^x

y = 3^(-3)

y = 1/(3^3)

y = 1/27

y = 0.037037

So we have one point (-3, 0.037037)
----------------

and when x = -2, then...

y = 3^x

y = 3^(-2)

y = 1/(3^2)

y = 1/9

y = 0.1111

So we have another point (-2, 0.1111)
----------------

and when x = -1, then...

y = 3^x

y = 3^(-1)

y = 1/(3^1)

y = 1/3

y = 0.3333

So we have another point (-1, 0.3333)
----------------

and when x = 0, then...

y = 3^x

y = 3^(0)

y = 1 ... Note: any number (except 0) to the zeroth power is 1.


So we have another point (0, 1)
----------------

and when x = 1, then...

y = 3^x

y = 3^(1)

y = 3 ... Note: any number to the first power is that same number


So we have another point (1, 3)
----------------


and when x = 2, then...

y = 3^x

y = 3^(2)

y = 9


So we have another point (2, 9)
----------------


and finally, when x = 3, then...

y = 3^x

y = 3^(3)

y = 27


So we have another point (3, 27)
----------------

Now we have the following points

(-3, 0.037037), (-2, 0.1111), (-1, 0.3333), (0, 1), (1, 3), (2, 9), (3, 27)


Now plot them on a coordinate grid:



Now draw a curve (as best you can) through all of the points to graph y = 3^x on the interval [-3, 3]


Note: take notice as the red graph of y = 3^x stops at (3,27) because it's at the end of the interval (the same applies to the other end of the graph).